OpenStudy (anonymous):

(log base a of 25)/(log base a of 3a)

OpenStudy (anonymous):

\[\log_{a}25 \div \log_{a} 3a\]

OpenStudy (anonymous):

Oops,sorry I forgot to add this \[\log_{a} 3 =0.477 ~and ~ \log_{a} 5 =0.699\]

OpenStudy (anonymous):

Note that \(\large \dfrac{\log_a 25}{\log_a(3a)} = \dfrac{\log_a(5^2)}{\log_a3 + \log_a a}=\ldots\) Can you take things from here? :-)

OpenStudy (anonymous):

Yup,but what if I were to apply product law for the numerator?

OpenStudy (anonymous):

You could do that as well by letting \(\large 25 = 5\cdot 5\). Then \(\large \log_a25 = \log_a(5\cdot 5) = \log_a5+\log_a5 = 2\log_a5 = \log_a(5^2)\). You can now see that both ways are equivalent. I hope this clarifies things! :-)

OpenStudy (anonymous):

Yes,it does,Thanks

OpenStudy (anonymous):

Um,could you give me the final answer just to check?

OpenStudy (anonymous):

You should end up with \(\large \dfrac{2\log_a 5}{\log_a3 + 1} = \dfrac{2\cdot 0.699}{0.477+1} \approx 0.946513 \)

OpenStudy (anonymous):

Ahh,now it makes sense.I completely forgot about the plus one.lol thanks